OpenStudy (anonymous):

Solve: ((2)/(3))^(2)÷((9)/(16))(3) show steps? Thank you.

OpenStudy (anonymous):

\[2/3^{2} \div 9/16(3) please help!\]

OpenStudy (sriram):

16(3) means that 16*3 in the denominator?

OpenStudy (anonymous):

no 9/16 times 3

OpenStudy (anonymous):

mathmaticalpro can u help me? lol

OpenStudy (anonymous):

\[\frac{(2/3)^2}{3(9/16)}\]?????

OpenStudy (anonymous):

(2/3)^2 div (9/16) (3) ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{(2/3)^2}{(9/16)(3)}\] \[\frac{(4/9)}{(27/16}\] \[(4/9)*(16/27)\] \[\frac{64}{243}\]

OpenStudy (anonymous):

Yesss! Correct, you are a math magician ! haha lol =D